OpenStudy (anonymous):

Solid sodium reacts violently with water, producing heat, hydrogen gas, and sodium hydroxide. How many molecules of hydrogen gas are formed when 48.7 g of sodium are added to water? 2Na + 2H2O --> 2NaOH + H2 Show your work.

OpenStudy (anonymous):

@ParthKohli ? @Mertsj ? Can you help?

OpenStudy (jfraser):

balanced reactions tell us the ratios between the pieces we use and the pieces we form, but it doesn't tell us MASS ratios, it tells us ratios of QUANTITY. To predict the amount of product we need to know the number of moles that the problem gives us. Can you turn 48.7g of Na into moles of Na?

OpenStudy (anonymous):

is it 2.12 moles of Na?

OpenStudy (anonymous):

@JFraser ?

OpenStudy (anonymous):

@dumbcow @ghazi @hba ? Can someone please help me???

OpenStudy (aaronq):

(1) convert to moles (2) use coefficients to determine how many moles of product you have eg. aA + bB -> cC moles of A/a = moles of C/c if you have 1 mole of A and want to find moles of C, then: moles of C=1*c/a (3) once you have the moles, covert to molecules knowing that 1 mole = avogadros number

OpenStudy (anonymous):

I'm lost...

OpenStudy (aaronq):

do step by step and post it, il check it

OpenStudy (anonymous):

is it converted to moles = 2.12 ?

OpenStudy (aaronq):

yes

OpenStudy (anonymous):

Ok so what do I do next? I'm sorry I'm really confused and I need to figure this out.

OpenStudy (aaronq):

use the 2nd step i outlined

OpenStudy (anonymous):

How do I use the coefficients to determine how many moles I have though??

OpenStudy (anonymous):

I'm just really really confused..

OpenStudy (aaronq):

look at the step i posted and follow what i did

OpenStudy (anonymous):

I know that but I'm obviously confused by what you did and I need help figuring it out....

OpenStudy (aaronq):

it's simple algebra, il give you another example: N2 + 3 H2 → 2 NH3 if you have 3 moles of N2 and you want to find how many moles of NH3 you can make then: \[\frac{ n _{N _{2}} }{ 1}=\frac{ n _{NH _{3}} }{ 2 }\] \[\frac{3}{ 1}=\frac{ n _{NH _{3}} }{ 2 }\] solve of moles of NH3

OpenStudy (aaronq):

solve for* not of

OpenStudy (anonymous):

I'm still confused

OpenStudy (aaronq):

well try it by following the example and post it, il rectify your mistakes

OpenStudy (anonymous):

I don't know how to try it! I don't understand it when it's not the actual numbers I'm using... My brain doesn't function that way.. Can you maybe just at least set it up for me and I can try and see if I can get it then??

OpenStudy (aaronq):

if i set it up I'm basically answering the question for you, so i'm not going to do that. but lets try this, what are the coefficients of each of the substances in question?

OpenStudy (anonymous):

2Na is 2, 2H2O is 2, 2NaOH is 2, ?

OpenStudy (aaronq):

well sure, but what are the ones they are asking about? hint: re-read the question

OpenStudy (anonymous):

Hydrogen gas and water?

OpenStudy (aaronq):

"How many molecules of hydrogen gas are formed when 48.7 g of sodium are added to water?" they are asking about hydrogen and sodium, not water

OpenStudy (aaronq):

now, what are the coefficients?

OpenStudy (anonymous):

2 and 1?

OpenStudy (anonymous):

Oh wait. hold on

OpenStudy (anonymous):

2 ,2, and 1?

OpenStudy (aaronq):

why are there 3 different ones? there is only 2 substances in question: hydrogen gas and sodium

OpenStudy (anonymous):

oh sorry. 2 and 1.

OpenStudy (aaronq):

which one is which

OpenStudy (anonymous):

1 of hydrogen gas and 2 of sodium

OpenStudy (aaronq):

okay, so divide each one by it's coefficient and set them equal to each other

OpenStudy (anonymous):

How do you divvide Na by 2?

OpenStudy (aaronq):

you divide the moles by the coefficient

OpenStudy (anonymous):

I don't understand. I've been working on this homework for 3 hours.. my brain hates me

OpenStudy (aaronq):

how many moles of Na did 48.7 g give you?

OpenStudy (anonymous):

2.12?

OpenStudy (aaronq):

so divide that by the coefficient

OpenStudy (anonymous):

1.06?

OpenStudy (aaronq):

set it up as a ratio though

OpenStudy (aaronq):

do the same for H2, and set them equal to each other, and POST IT

OpenStudy (anonymous):

How do I set it as a ratio?

OpenStudy (aaronq):

a ratio is a fraction

OpenStudy (anonymous):

oh. ok

OpenStudy (aaronq):

0.5=1/2

OpenStudy (anonymous):

53/50 = 121/5 .....? I did that totally wrong didn't i

OpenStudy (aaronq):

okay, you're overcomplicating this, you're merely supposed to write: \[\frac{ moles of Na }{ coefficient }\]

OpenStudy (anonymous):

Omg. See, this is why I suck at school.

OpenStudy (aaronq):

it's not that you "suck", you just have to think about what you're trying to do and really think and anaylize it

OpenStudy (anonymous):

ok so would it be: 2.12 mol Na/2 = 24.2 mol H2/1 ?

OpenStudy (aaronq):

well yes and no. you don't know how many moles of H2 you have so that is your unknown, so 2.12/2=x/1 x=moles of H2

OpenStudy (anonymous):

Ok so is x=mol of H2 the answer?

OpenStudy (aaronq):

no, they're asking for the amount of molecules, so whatever you get for the moles of H2, multiply by avogadros number

OpenStudy (anonymous):

......... I'm lost again

OpenStudy (aaronq):

okay how many moles of H2 gas do you have?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

NO, 24.2

OpenStudy (anonymous):

NO, x

OpenStudy (aaronq):

x is not a number, you're solving for x from the equation above.

OpenStudy (aaronq):

2.12/2=x/1 what is the value of x?

OpenStudy (anonymous):

24.2?

OpenStudy (aaronq):

how did you get that?

OpenStudy (anonymous):

Uh. I don't remember..

OpenStudy (aaronq):

well do it again and find it because that previous answer is incorrect

OpenStudy (anonymous):

You're no help at all.

OpenStudy (aaronq):

lol okay then

OpenStudy (aaronq):

maybe it's you that is dense, i tried simplifying things to a grade 2 level, clearly you don't have the brain for science, ingrate.

OpenStudy (notstar):

Just for the record guys and gals reading this that got to this point, Once you get 1.06 for 2.12/2=x/1 which is pretty much 2.12 divided by 2. You should multiply it by Avogadro's number to get your final answer. Right? @aaronq

OpenStudy (aaronq):

yes